A generic iron preparation has a concentration of 20 mg/ml. Calculate the number of milliliters required to administer 200 mg/day for 180 days.
Answer #1

available medication 20mg/ml

required dose 200mg/day

how much ml will be required?

20 mg will be in 1ml

then for 180 day require ml=180x10=1800 ml

1 mg will be from=1/20 ml

200mg from = (1/20)X200=10ml

hence 10 ml of this medication will be required for the fulfillment of the order.

for 180 days required ml=180X10=1800ml